Using Rhodes Corporation’s financial statements (shown after part f), answer the following questions.

a. What is the net operating profit after taxes (NOPAT) for 2018?

b. What are the amounts of net operating working capital for both years?

c. What are the amounts of total net operating capital for both years?

d. What is the free cash flow for 2018?

e. What is the ROIC for 2018?

f. How much of the FCF did Rhodes use for each of the following purposes:

after-tax interest, net debt repayments, dividends, net stock repurchases,

and net purchases of short-term investments? (Hint: Remember that a net

use can be negative.)

Rhodes Corporation: Income Statements for Year Ending

December 31 (Millions of Dollars)

2018 2017

Sales $11,000 $10,000

Operating costs excluding depreciation 9,360 8,500

Depreciation and amortization 380 360

Earnings before interest and taxes $ 1,260 $ 1,140

Less interest 120 100

Pre-tax income $ 1,140 $ 1,040

Taxes (40%) 456 416

Net income available to common

stockholders $ 684 $ 624

Common dividends $ 220 $ 200

Continued...

Rhodes Corporation: Balance Sheets as of December 31

(Millions of Dollars)

2018 2017

Assets

Cash $ 550 $ 500

Short-term investments 110 100

Accounts receivable 2,750 2,500

Inventories 1,650 1,500

Total current assets $5,060 $4,600

Net plant and equipment 3,850 3,500

Total assets $ 8,910 $8,100

Liabilities and Equity

Accounts payable $ 1,100 $1,000

Accruals 550 500

Notes payable 384 200

Total current liabilities $2,034 $1,700

Long-term debt 1,100 1,000

Total liabilities $3,134 $2,700

Common stock 4,312 4,400

Retained earnings 1,464 1,000

Total common equity $5,776 $5,400

Total liabilities and equity $8,910 $8,100
Answer #1

we are supposed to solve 4 subparts when multiple posted

a)NOPAT=EBIT*(1-tax)
=1260*(1-40%)=756
b)net operating working capital =(cash + accounts receivable + inventories) – (accounts payable + accruals
2018=(550+2750+1650)-(1100+550)=3300
2017=(500+2500+1500)-(1000+500)=3000
c)total net operating capital =NOWC+operating long term assets
2018=3300+3850=7150
2017=3000+3500=6500
d)FCF= NOPAT + Net investment in operational capital
=756+(7150-6500)=1406